Remsen 101 Seating Layout File rem101, Revision B, jmb 11/98 Johns Hopkins University

There are 192 seats. All of them are right-handed. There are a few extra movable chairs at the back. With alternate seating, there is space for 102 people (or 101 if one allows for the broken seats). Every third seat is labeled. There is no row I.

* Several seats in row J have broken tables.

EVEN NUMBERS ARE ON THIS SIDE (WEST) ODD NUMBERS ARE ON THIS SIDE (EAST)

8* 6 4* 2* J  J 1 3* 5 7

22 20 18 16 14 12 10 8 6 4 2 H  H 1 3 5 7 9 11 13 15 17 19 21

26 24 22 20 18 16 14 12 10 8 6 4 2 G 1 3 5 7 9 11 13 15 17 19 21 23 25

26 24 22 20 18 16 14 12 10 8 6 4 2 F 1 3 5 7 9 11 13 15 17 19 21 23 25

26 24 22 20 18 16 14 12 10 8 6 4 2 E 1 3 5 7 9 11 13 15 17 19 21 23 25

24 22 20 18 16 14 12 10 8 6 4 2 D 1 3 5 7 9 11 13 15 17 19 21 23

22 20 18 16 14 12 10 8 6 4 2 C 1 3 5 7 9 11 13 15 17 19 21

20 18 16 14 12 10 8 6 4 2 B 1 3 5 7 9 11 13 15 17 19

18 16 14 12 10 8 6 4 2 A 1 3 5 7 9 11 13 15 17

BENCH